OpenStudy (anonymous):

helpppp! a recent survey reveals that of the 250 poodle owners interviewed \,93% of them regularly groom their dogs.if there are 5,000 registered poodle owners in the city,how many owners are expected to regularly groom their dogs?

OpenStudy (anonymous):

is that 93%?

OpenStudy (anonymous):

i believe so thats what i got

OpenStudy (anonymous):

Since you still expect 93% of them to do it, youll find that 93% of these 5000 are likely to groom their poodles. So thats 0.93 * 5000 = 93 * 50 = 4650

OpenStudy (anonymous):

thank you!